OpenStudy (anonymous):

Which of the following statements is true? Choose one answer. a. An exponential function is the inverse of a logarithmic function. b. The base of the logarithmic function f(x) = log5x is unknown. c. The range of the logarithmic function f(x) = log5x is all real numbers less than zero. d. The domain of the logarithmic function f(x) = log5x is all real numbers.

OpenStudy (curry):

A

OpenStudy (curry):

to my knowledge

OpenStudy (curry):

imran check it for me

OpenStudy (anonymous):

a

OpenStudy (anonymous):

y=e^x change variable x= e^y y ln e = ln x y= ln x